OpenStudy (shadow13):

Which relation is a function? {(1, 2); (2, 3); (3, 4); (4, 5)} {(1, 2); (1, 3); (1, 4); (1, 5)} {(1, 2); (2, 3); (3, 4); (2, 5)} {(1, 2); (3, 2); (3, 3); (4, 2)}

OpenStudy (shadow13):

@endlesslight254

OpenStudy (shadow13):

i think it is A

OpenStudy (anonymous):

i think so to but im not sure .3.

OpenStudy (shadow13):

ok

OpenStudy (shadow13):

What is the range of the relation? {(1, 2), (2, 4), (3, 1), (4, 3)} {1, 2, 3, 4} {1, 2, 3, 4, 6} {2, 4, 6} {2, 4} i think it is d @endlesslight254

OpenStudy (anonymous):

d or b

OpenStudy (shadow13):

ok what does it mean when its asking to find the range of a relation i have a problem on that and its hard to understand it

OpenStudy (anonymous):

let me look it up im bad at these problems lol

OpenStudy (shadow13):

ok

OpenStudy (anonymous):

i hope this helps love <3 State the domain and range of the following relation. Is the relation a function? {(2, –3), (4, 6), (3, –1), (6, 6), (2, 3)} The above list of points, being a relationship between certain x's and certain y's, is a relation. The domain is all the x-values, and the range is all the y-values. To give the domain and the range, I just list the values without duplication: domain: {2, 3, 4, 6} range: {–3, –1, 3, 6}

OpenStudy (shadow13):

ok now that makes since then it is b

OpenStudy (anonymous):

yay ^u^

OpenStudy (anonymous):

want the website i found?

OpenStudy (shadow13):

no im good thanks though

OpenStudy (anonymous):

kk ^u^

OpenStudy (shadow13):

For the function f(x) = 2 – 3x, find f(–4). –10 –5 12 14 other people are saying its 12 but i get -5

OpenStudy (anonymous):

ohhh im good at these!

OpenStudy (shadow13):

yay

OpenStudy (anonymous):

F(-4)=2-3(-4) f(-4)=2--12 f(-4)=2+12 f(-4)=14 the answer is 14

OpenStudy (shadow13):

oh ok

OpenStudy (anonymous):

^u^

OpenStudy (shadow13):

i suck at algebra

OpenStudy (shadow13):

What is the turning point of the graph of f(x) = |x – 4| ? (0, 4) (0, –4) (4, 0) (−4, 0)

OpenStudy (anonymous):

its ok i wasnt good at it either

OpenStudy (shadow13):

i think it is b

OpenStudy (anonymous):

x - 4 = 0 x = 4 f(4) = |x - 4| turning point : (4, 0)

OpenStudy (anonymous):

im surprised i even know this stuff

OpenStudy (shadow13):

Which is the graph of the function g(x) = |x – 3|? (Points : 4) Which is the graph of the function g(x) = |x – 3|? (Points : 4) Which is the graph of the function g(x) = |x – 3|? (Points : 4) Which is the graph of the function g(x) = |x – 3|? (Points : 4) Which is the graph of the function g(x) = |x – 3|? it needs to be a gragh and i think it is d but not sure

OpenStudy (shadow13):

oops tomany

OpenStudy (anonymous):

XD

OpenStudy (shadow13):

can u help though there v graphed and its up or down

OpenStudy (anonymous):

i can try could u post the graph??

OpenStudy (shadow13):

it wont let me

OpenStudy (anonymous):

thats weird 'n'

OpenStudy (shadow13):

thats the graphs

OpenStudy (shadow13):

thats d

OpenStudy (shadow13):

heres c

OpenStudy (shadow13):

b

OpenStudy (shadow13):

a

OpenStudy (anonymous):

which one at the top is the question?

OpenStudy (shadow13):

Which is the graph of the function g(x) = |x – 3|? this is the question for those answers

OpenStudy (anonymous):

im bad at this question put what ever u think it is love ^u^

OpenStudy (shadow13):

ok

OpenStudy (shadow13):

Which is the direct linear variation equation for the relationship? y varies directly with x and y = 10 when x = 2. y = 2x + 6 y = x + 8 y = x – 8 y = 5x

OpenStudy (shadow13):

i think it is b

OpenStudy (anonymous):

its a or b ^u^

OpenStudy (shadow13):

its a

OpenStudy (shadow13):

but other people say it is d

OpenStudy (anonymous):

then its probably d cause of the 10 but then again it also could be a

OpenStudy (shadow13):

yeah thats true ill try d

OpenStudy (anonymous):

alright ^u^

OpenStudy (shadow13):

Which relationship shows a quadratic variation? A X: 2,3,4,5 f(x): 1,4,9,16 b. x: 1,2,3,4 f(x):2,8,18,32 c. x:1,2.3,4 f(x): 2,5,10,17 d. x: 1,2,3,4 f(x): 4,3,2,1

OpenStudy (shadow13):

i think it is a

OpenStudy (anonymous):

yea it must be a :3

OpenStudy (shadow13):

ok

OpenStudy (shadow13):

Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 36 when x = 3. y = x2 y = x2 y = 4x2 y = 12x2 think it is B

OpenStudy (anonymous):

its b or c

OpenStudy (anonymous):

ur bro got a account and now asking me to help him XD

OpenStudy (shadow13):

dont do it

OpenStudy (anonymous):

im not gonna help him when im helping u <3

OpenStudy (shadow13):

Which relationship shows an inverse variation? A.x. 2,3,4,5 f(x):1,4,9,16 B. x.1,223,4 f(x): 2,8,18,32 C.x:1,2,3,4 f(x):4,3,2,1 D.x: 2,4,6,8 f(x): 12,6,4,3 i dont know

OpenStudy (anonymous):

idk either O.O

OpenStudy (shadow13):

if i took a wild guess i would pic b

OpenStudy (anonymous):

same lol

OpenStudy (anonymous):

are u doing a math test or a assignment?

OpenStudy (shadow13):

math test i have two more

OpenStudy (anonymous):

i hope u pass :o

OpenStudy (shadow13):

y varies inversely with x and y = 8 when x = 2. Which is the inverse variation equation for the relationship? (Points : 4) y = 16x y=16/x y = 2x y = x + 2 i think it is c

OpenStudy (shadow13):

i hope so too

OpenStudy (anonymous):

its c ^u^

OpenStudy (shadow13):

yay i did it

OpenStudy (shadow13):

If f(x) varies directly with x and f(x) = 50 when x = –10, then what is f(x) when x = 15 –150 –75 –30 –5 i think it is B

OpenStudy (anonymous):

yay im glad u got that question right :D

OpenStudy (shadow13):

why

OpenStudy (anonymous):

cause that means ur smart silly ^u^

OpenStudy (shadow13):

o ok

OpenStudy (anonymous):

^u^

OpenStudy (shadow13):

i got more quiez and exams and assinments

OpenStudy (shadow13):

ill ask for ur help when i need it

OpenStudy (anonymous):

and its either a or b

OpenStudy (shadow13):

ok be back soon for help

OpenStudy (anonymous):

alright ill miss u <3

OpenStudy (shadow13):

same ily

OpenStudy (anonymous):

ilyt